# NACHOS: Neural Architecture Search for Hardware Constrained Early Exit Neural Networks

### v1

@@ V-C Analysis of the results
The experimental results of NACHOS on the three different datasets are presented in Table I, Table II, and Table III. The column NAS specifies if the network is designed by a NAS procedure (yes) or not (no). The column B specifies the number of EECs (including the final exit). The column $F_{M}(\widetilde{x})$ shows the number of MACs in EE inference. The column $F_{A}(\widetilde{x})$ shows the top1 accuracy in EE inference. The column $F_{A}(x)$ shows the top1 accuracy in inference by using only the backbone of the EENN. The column $\theta$ shows the thresholds of the EECs (apart from the last classifier not having a thresholding). The column U shows the exit ratios $U_{i}$ of the EECs. The column Confidence shows the technique used to estimate the confidence in the prediction.
@@
Additionally, the computational costs $\gamma_{i}$, i.e., the amounts of MACs to process an input at the $i^{th}$ exit of an EENN, of the models presented in the aforementioned tables are reported in Table IV with one column for each dataset. We highlight that the costs of EEAlexnet and EEResnet20 are the same for each dataset since they refer to the same model. For this reason, the value is set to Same.
Comparing NACHOS with the other state-of-the-art EENNs we show that NACHOS is effective in selecting a competitive network in terms of best trade-off between the two objectives of the search. In particular, NACHOS is able to find an EENN with similar computational demand to EDANAS with similar accuracy on CINIC-10 while improving the accuracy on CIFAR-10 and SVHN. Moreover, NACHOS is competitive with the expert-designed models on CIFAR-10 and CINIC-10, while it performs worse on SVHN. We highlight that the computational demand of the expert-designed models is significantly higher with respect to NACHOS.
Additionally, the Expected Calibration Error (ECE) scores for the experimental campaign of NACHOS on CIFAR-10 dataset are presented in Table V. ECE is a popular metric to measure the calibration of the EECs [54]. The calibration is an important aspect of EENNs related to their capability of estimating properly the confidences on input samples.
Table V shows that the ECE scores of the NAS model are good pointing out that well-calibrated EENNs are found by the NACHOS method. However, we expect calibration to play a major role in NACHOS when considering networks with higher computational demands since, for instance, deeper networks tend to be more miscalibrated [55] (e.g., with Resnet-based OFA). This aspect will be addressed in future works of the paper.
@@ V-D Ablation studies
TABLE VII: Results and computational costs ($\gamma$) for the ablation studies of NACHOS on CIFAR-10 dataset. Column labels are explained in Table VI.
Model NAS B $\gamma$ $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\theta$ U (%) Confidence
NACHOS yes 2 [2.41, 5.95] 2.44 5.95 72.86 71.41 [0.2] [99.0, 1.0] Cumulative
  NACHOS_NC yes 3 [1.74, 2.95, 6.48] 2.81 6.48 72.18 71.09 [0.7, 0.1] [11.0, 89.0, 0.0] Cumulative
  NACHOS_NP yes 2 [2.30, 6.66] 2.36 6.66 71.72 68.65 [0.2, 0.1] [99.0, 1.0] Cumulative
TABLE VIII: Results and computational costs ($\gamma$) for the ablation studies of NACHOS on CINIC-10 dataset. Column labels explained in Table VI.
Model NAS B $\gamma$ $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\theta$ U (%) Confidence
NACHOS yes 2 [2.28, 5.97] 2.42 5.97 60.77 59.65 [0.2,0.1] [96.0, 4.0] Cumulative
  NACHOS_NC yes 3 [3.45, 3.65, 6.98] 3.82 6.98 59.86 59.66 [0.7,0.5, 0.5] [5.91, 88.68, 5.41] Cumulative
  NACHOS_NP yes 2 [2.10, 4.83] 2.33 4.83 60.44 59.81 [0.3, 0.1] [92.0, 8.0] Cumulative
TABLE IX: Results and computational costs ($\gamma$) for the ablation studies of NACHOS on SVHN dataset. Column labels explained in Table VI.
Model NAS B $\gamma$ $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\theta$ U (%) Confidence
NACHOS yes 3 [1.17, 1.88, 4.61] 1.49 4.61 79.98 88.71 [0.3,0.0,0.1] [56.0, 44.0, 0.0] Cumulative
  NACHOS_NC yes 3 [1.27, 1.99, 4.71] 1.84 4.71 84.87 88.12 [0.7, 0.1, 0.1] [20.0, 80.0, 0.0] Cumulative
  NACHOS_NP yes 3 [1.18, 1.96, 4.68] 1.46 4.68 77.18 89.31 [0.3, 0.0, 0.1] [64.4, 35.6, 0.0] Cumulative
@@
Figure 4: Comparison between the distribution of confidences of NACHOS models trained by applying different regularizations.
@@
The results and the computational costs of the models NACHOS_NC and NACHOS_NP found on three datasets are reported in Table VII, Table VIII, and Table IX.
@@
In the NACHOS experiments, comparing NACHOS with NACHOS_NC, we aim at testing the effectiveness of the constraints in selecting an EENN satisfying the given constraints. The results show that the constraints are satisfied in NACHOS while are not satisfied in NACHOS_NC. Additionally, Fig. 5 points out that NACHOS with constraints is able to reach faster the region of solutions satisfying the constraints, delimited by the red horizontal line and the purple vertical line in the figure, with respect to the unconstrained counterpart. Finally, we can argue that a NAS by enforcing constraints is able to find an admissible solution with a significant lower number of iterations. Indeed, NACHOS and EDANAS models are found by NACHOS in 10 iterations and by EDANAS in 30 iterations, respectively.
Figure 5: Comparison between the Pareto fronts of NACHOS enforcing and not enforcing constraints in the search on CIFAR-10 respectively.
Fig. 6 shows the utilization of the second EEC of the NACHOS_NC on CIFAR-10 varying the thresholds of the first two EECS. This visualizes how the thresholds affect the behaviour of the Cumulative method in triggering the early exit in the EECs.
Figure 6: Heatmap of the utilization of the second EEC of the NACHOS_NC varying the thresholds of the first two EECs.
In the NACHOS experiments, comparing NACHOS with NACHOS_NP we aim at testing the effectiveness of the $L_{peak}$ regularization in improving the accuracy of the EENN. An increase in accuracy is seen on CIFAR-10 and SVHN while the accuracy is approximately the same on CINIC-10.
@@
Lastly, we report in Fig. 4 the distributions of the confidences in NACHOS, NACHOS_NP, and NACHOS_NC models. We can point out that the $L_{peak}$ is effective in aligning the confidences towards the operating point of the EECs. This is evident in the decrease of the number of samples for very high confidences for both EECs and the corresponding increase of samples towards the operating point in the NACHOS model with respect to the NACHOS_NP model. Moreover, the $L_{cost}$ is effective in favoring the training of the early fraction of the EENN. This is seen in the higher number of samples for the higher confidences in the NACHOS model in the first EEC with respect to the NACHOS_NC model.
@@ VI Conclusions
In this work, we have presented Neural Architecture Search for Hardware COnstrained Early Exit Neural Networks (NACHOS), the first NAS framework for the design of optimal EENNs satisfying a constraint on the number of Multiply and Accumulate (MAC) operations performed by the EENNs at inference time. The results show that NACHOS is able to select a set of Pareto Optimal Solutions competitive with the other state-of-the-art EENNs. Additionally, we investigated the effectiveness of the novel regularizations introduced in the training of the EENNs proposed in NACHOS. Future works will examine different OFAs, early-exit schemes, optimization techniques, application scenarios, and the combination of NACHOS with other AutoML techniques.

### v2

@@ V-C Analysis of the results
The experimental results of NACHOS are presented in Table II. The column NAS specifies if the network is designed by a NAS procedure (yes) or not (no). The column B specifies the number of EECs (including the final exit). The column $F_{M}(\widetilde{x})$ shows the number of MACs in EE inference  [^2] The number of Floating Point Operations (FLOPS), another popular metric regarding inference time, can be approximately computed by the number of MACs as $\text{FLOPS}\approx 2\times\text{MACs}$. The column $F_{A}(\widetilde{x})$ shows the top1 accuracy in EE inference. The column $F_{A}(x)$ shows the top1 accuracy in inference by using only the backbone of the EENN. The column $\epsilon$ shows the thresholds of the EECs (apart from the last classifier not having a thresholding). The column Confidence shows the technique used to estimate the confidence in the prediction.
Additionally, the computational costs $\gamma_{i}$, i.e., the amounts of MACs to process an input at the $i^{th}$ exit of an EENN, and the utilizations $U_{i}$, i.e., the exit ratios of the EECs, of the models presented in the aforementioned tables are reported in Table III. We highlight that the costs of EEAlexnet and EEResnet20 are the same for SVHN, CIFAR-10, and CINIC-10 since they refer to the same model.
@@
Comparing NACHOS with the other state-of-the-art EENNs we show that NACHOS is effective in selecting a competitive network in terms of best trade-off between the two objectives of the search. In particular, NACHOS is able to find an EENN with similar computational demand to EDANAS with similar accuracy on CINIC-10 while improving the accuracy on CIFAR-10, SVHN, and Imagenette. Moreover, NACHOS is competitive with the expert-designed models on CIFAR-10, CINIC-10, and Imagenette while it performs worse on SVHN. We highlight that the computational demand of the expert-designed models is significantly higher with respect to NACHOS.
Additionally, the Expected Calibration Error (ECE) scores for the experimental campaign of NACHOS on CIFAR-10 dataset are presented in Table IV. ECE is a popular metric to measure the calibration of the EECs [61]. The calibration is an important aspect of EENNs related to their capability of estimating properly the confidences on input samples.
Table IV shows that the ECE scores of the NAS model are good pointing out that well-calibrated EENNs are found by the NACHOS method. However, we expect calibration to play a major role in NACHOS when considering networks with higher computational demands since, for instance, deeper networks tend to be more miscalibrated [62] (e.g., with Resnet-based OFA). This aspect will be addressed in future works of the paper.
@@ V-D Ablation studies
Figure 5: Comparison between the distribution of confidences of NACHOS models trained by applying different regularization while setting the same seed.
@@
The results of the models NACHOS, NACHOS_NC and NACHOS_NP found on the four datasets are reported in Table V. Additionally, the computational costs and the utilizations of the EECs of the models NACHOS, NACHOS_NC, and NACHOS_NP found on the four datasets are reported in Table VI.
@@
In the NACHOS experiments, comparing NACHOS with NACHOS_NC, we aim at testing the effectiveness of the $L_{cost}$ regularization in selecting an EENN satisfying the given constraints. The results show that the constraints are satisfied in NACHOS while are not satisfied in NACHOS_NC, Additionally, the model found shows even better accuracy on CIFAR-10 and SVHN. Additionally, Fig. 4 points out that NACHOS by using constraints and the $L_{cost}$ regularization is able to reach faster the region of solutions satisfying the constraints, delimited by the red horizontal line and the purple vertical line in the figure, with respect to the unconstrained counterpart. Finally, we can argue that a NAS by enforcing constraints is able to find an admissible solution with a significant lower number of iterations. Indeed, NACHOS and EDANAS models are found by NACHOS in 10 iterations and by EDANAS in 30 iterations, respectively. Fig. 4 shows the utilization of the second EEC of a model with three exits (B=3) found by NACHOS_NC on CIFAR-10 varying the thresholds of the first two EECS. This visualizes how the thresholds affect the behaviour of the Cumulative method in triggering the early exit in the EECs. In the NACHOS experiments, comparing NACHOS with NACHOS_NP we aim at testing the effectiveness of the $L_{peak}$ regularization in improving the accuracy of the EENN. An increase in accuracy is seen on CIFAR-10, SVHN, and Imagenette while the accuracy is approximately the same on CINIC-10 and maintains a similar computational complexity.
@@
Lastly, we report in Fig. 5 the distributions of the confidences in NACHOS, NACHOS_NP, and NACHOS_NC models obtained with the same seed. We can point out that the $L_{peak}$ is effective in aligning the confidences towards the operating point of the EECs. This is evident in the decrease of the number of samples for very high confidences for both EECs and the corresponding increase of samples towards the operating point in the NACHOS model with respect to the NACHOS_NP model. Moreover, the $L_{cost}$ is effective in favoring the training of the early fraction of the EENN. This is seen in the higher number of samples for the higher confidences in the NACHOS model in the first EEC with respect to the NACHOS_NC model.
V-E Statistical tests
To validate the statistical significance of our results, we performed the Friedman statistical test by using $F_{M}(\tilde{x})$ and $F_{A}(\tilde{x})$ of the multiple runs over five seeds of all our tested methods, i.e., NACHOS, NACHOS_NP, NACHOS_NC, and EDANAS, for each dataset and reported results in Table VII. The results shows that there are statistic differences among the methods, hence the competitive performance of NACHOS is statistically validated.
TABLE VII: Friedman test results for each dataset and objective. The higher $\chi^{2}$, the more different the results of the methods based on their rankings and the smaller the p-value (typically $<0.05$), the more statistically significant the $\chi^{2}$ statistic.
Dataset & Objective Chi-squared ($\chi^{2}$) p-value
SVHN ($F_{M}(\tilde{x})$) 9.9600 0.0189
SVHN ($F_{A}(\tilde{x})$) 15.0000 0.0018
CIFAR-10 ($F_{M}(\tilde{x})$) 14.0400 0.0029
CIFAR-10 ($F_{A}(\tilde{x})$) 13.5600 0.0036
CINIC-10 ($F_{M}(\tilde{x})$) 10.5625 0.0143
CINIC-10 ($F_{A}(\tilde{x})$) 13.0408 0.0045
Imagenette ($F_{M}(\tilde{x})$) 15.0000 0.0018
Imagenette ($F_{A}(\tilde{x})$) 15.0000 0.0018
@@ VI Conclusions
In this work, we have presented Neural Architecture Search for Hardware Constrained Early Exit Neural Networks (NACHOS), the first NAS framework for the design of optimal EENNs satisfying a constraint on the number of Multiply and Accumulate (MAC) operations performed by the EENNs at inference time. The results show that NACHOS is able to select a set of Pareto Optimal Solutions competitive with the other state-of-the-art EENNs. Additionally, we investigated the effectiveness of the novel regularizations introduced in the training of the EENNs proposed in NACHOS. Future works will examine different OFAs, early-exit schemes, optimization techniques, application scenarios, and the combination of NACHOS with other AutoML techniques. In particular, we stress that one limitation of the current research is that it primarily focuses on mobile devices, optimizing neural networks with constraints in the range of millions of parameters and MACs. Future research directions could focus on further optimization of neural networks by incorporating hardware accelerator architectures and compiler mappings directly into the design space. This could lead to the development of models that are efficient enough to be deployed on MCUs and in general to scenarios with very severe constraints.